• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 14:25:17 on June 3,2010

    Entry number 324270

    keyword=prep work for data taking

    Had some problems with adaql2 (separate halog).

    Got adaql2 to work and started L-HRS and R-HRS runs with
    pulser trigger. Ahmed Zafar had done some cosmics checkout
    of the detectors yesterday and e-mailed that VDCs and S0
    are good to go. I didn't see an halog about that.

    Turned off VDC HV just now.

    Noted the HRS magnets are all on.

    The target is being cooled down now.